# MMLA: Multi-Environment, Multi-Species, Low-Altitude Drone Dataset

### v2

@@ 1 Introduction
To address this gap, we present a novel multi-environment, multi-species, low-altitude drone dataset (MMLA), the first low-altitude dataset explicitly designed for developing robust computer vision models for autonomous wildlife-tracking drones. Our dataset spans three distinct locations (Ol Pejeta Conservancy and Mpala Research Centre in Kenya and The Wilds Conservation Center in Ohio, USA) and six species: plains zebras, Grevy’s zebras, reticulated and Masai giraffes, Persian onagers, and African wild dogs. MMLA contains 155K frames (1 hr, 26 min), with manually annotated bounding boxes around each animal optimized for training high-precision detection models.
@@ 3 MMLA Dataset
The MMLA dataset [^1] https://huggingface.co/collections/imageomics/mmla-project contains annotated drone collected for wildlife monitoring and conservation research. The data was collected at three locations: Ol Pejeta Conservancy and Mpala Research Centre, both in Laikipia, Kenya, and The Wilds Conservation Center, in Ohio, USA. The dataset contains six species: Plains zebra (Equus quagga), Grevy’s zebra (Equus grevyi), Persian onager (Equus hemionus onager), Reticulated giraffe (Giraffa reticulata), Masai giraffe (Giraffa camelopardalis tippelskirchi), and African Wild Dog (Lycaon pictus). The variety of species and environments in the MMLA dataset enables the evaluation of model generalizability in detecting animals across different settings. The dataset includes 155,074 video frames, captured at 30 frames per second, and contains a variety of habitats (details in App. 9 Tab. 4). Each video frame has COCO format annotations (bounding boxes with normalized coordinates). The dataset contains 811K total annotations, which were completed manually by a team of experts using CVAT [7] and kabr-tools [19].
The Ol Pejeta dataset contains 29,268 frames of Plains zebras at Ol Pejeta Conservancy in Kenya (Jan-Feb 2025), collected using the autonomous WildWing system [18] across seven video sessions during the 2025 WildDrone Hackathon [36]. The Mpala dataset contains 104,062 frames of reticulated giraffes, Grevy’s and Plains zebras at Mpala Research Center in Kenya (Jan 2023), collected manually using a DJI Air 2S drone across five sessions [15]. The Wilds dataset contains 21,744 frames of African wild dogs, Persian onagers, reticulated and Masai giraffes and Grevy’s zebras at The Wilds Conservation Center in Ohio (Apr-Jul 2024), collected using both DJI Mini and Parrot Anafi drones, with with a combination of autonomous WildWing system [16] and manual collection.
@@
YOLOv11m Finetuned on MMLA 74.39 83.43 81.39 81.76 82.40
@@ 4.2 Fine-tuned YOLOv11m
We trained the model for 50 epochs (approximately 2 hours) on the Ohio Supercomputer Cluster, using two Tesla V100-PCIE-16GB GPUs. The hyperparameters and optimizers were left as the default for YOLO fine-tuning using scripts provided in the ultralytics library [14]. We provide scripts to reproduce our results in our GitHub repository and provide the fine-tuned weights [^3] https://huggingface.co/imageomics/mmla. Additional training details may be found in App. 11.
@@ 7 Conclusion
Low-altitude drone studies offer a valuable middle ground between the detailed imagery of camera traps and the broad coverage of high-altitude surveys, enabling new behavioral and movement studies. Autonomous drones combined with robust detection models can transform wildlife monitoring from a static, pre-planned process to an adaptive system that responds to animal movements in real-time [18]. Deploying such systems involves important tradeoffs for ecologists. Lower flight altitudes provide better resolution but may increase animal disturbance, potentially altering natural behaviors [1, 30]. The balance between human and automated analysis also presents challenges [17]. Manual analysis by trained biologists remains the reliability standard but limits study scale, while automated detection can dramatically increase processing speed but requires domain-specific training to achieve necessary reliability.
The dramatic improvements achieved through fine-tuning suggest that with appropriate dataset development, AI-assisted wildlife monitoring can achieve the reliability needed for scientific applications while significantly increasing wildlife studies’ temporal and spatial coverage. By contributing both the MMLA dataset and fine-tuned model weights to the research community, we hope to accelerate the development of robust computer vision solutions for wildlife conservation and establish a foundation for the next generation of intelligent, autonomous wildlife monitoring systems.
@@ 8 Acknowledgments
We want to thank Samuel Mutisya and William Njoroge at Ol Pejeta Conservancy for their support in facilitating the WildDrone 2025 Hackathon. This work was supported by WildDroneEU (EU Horizon Europe MSCA grant 101071224), the Imageomics Institute (NSF HDR Award 2118240), and ICICLE (NSF grant OAC-2112606).

### v3

@@ 1 Introduction
To address this gap, we present a novel multi-environment, multi-species, low-altitude (MMLA) drone dataset  [^1] MMLA - an Imageomics HuggingFace Collection, the first low-altitude dataset explicitly designed for developing robust computer vision models for autonomous wildlife-tracking drones. Our dataset spans three distinct locations (Ol Pejeta Conservancy and Mpala Research Centre in Kenya and The Wilds Conservation Center in Ohio, USA) and six species: plains zebras, Grevy’s zebras, reticulated and Masai giraffes, Persian onagers, and African wild dogs. MMLA contains 155K frames (1 hr, 26 min), with manually annotated bounding boxes around each animal optimized for training high-precision detection models.
@@ 3 MMLA Dataset
The MMLA dataset contains annotated drone collected for wildlife monitoring and conservation research. The data was collected at three locations: Ol Pejeta Conservancy and Mpala Research Centre, both in Laikipia, Kenya, and The Wilds Conservation Center, in Ohio, USA. The dataset contains six species: Plains zebra (Equus quagga), Grevy’s zebra (Equus grevyi), Persian onager (Equus hemionus onager), Reticulated giraffe (Giraffa reticulata), Masai giraffe (Giraffa camelopardalis tippelskirchi), and African wild dog (Lycaon pictus). The variety of species and environments in the MMLA dataset enables the evaluation of model generalizability in detecting animals across different settings. The dataset includes 155,074 video frames, captured at 30 frames per second, and contains a variety of habitats (details in App. 9 Tab. 4). Each video frame has COCO format annotations (bounding boxes with normalized coordinates). The dataset contains 811K total annotations, which were completed manually by a team of experts using CVAT [7] and kabr-tools [21].
The Ol Pejeta dataset  [^2] https://huggingface.co/datasets/imageomics/mmla_opc contains 29,268 frames of Plains zebras at Ol Pejeta Conservancy in Kenya (Jan-Feb 2025) [23]. The videos were collected using the semi-autonomous WildWing system [26] across seven video sessions during the 2025 WildDrone Hackathon [43]. The Mpala dataset [^3] https://huggingface.co/datasets/imageomics/mmla_mpala contains 104,062 frames of reticulated giraffes, Grevy’s and Plains zebras at Mpala Research Center in Kenya (Jan 2023) [22]. The Mpala data was collected manually using a DJI Air 2S drone across five sessions [20, 19]. The Wilds dataset [^4] https://huggingface.co/datasets/imageomics/mmla_wilds contains 21,744 frames of African wild dogs, Persian onagers, reticulated and Masai giraffe, and Grevy’s zebras [14, 18]. This data was collected at The Wilds in Ohio (Apr-Jul), using both DJI Mini and Parrot Anafi drones, with a combination of manually piloted flights and the semi-autonomous WildWing system [26, 25].
@@
YOLOv11m Finetuned on MMLA [17] 74.39 83.43 81.39 81.76 82.40
@@ 4.2 Fine-tuned YOLOv11m
We trained the model for 50 epochs (approximately 2 hours) on the Ohio Supercomputer Cluster, using two Tesla V100-PCIE-16GB GPUs. The hyperparameters and optimizers were left as the default for YOLO fine-tuning using scripts provided in the ultralytics library [15]. We provide scripts to reproduce our results in our GitHub repository and provide the fine-tuned weights  [^6] https://huggingface.co/imageomics/mmla [17]. Additional training details may be found in App. 11.
@@ 7 Conclusion
Low-altitude drone studies offer a valuable middle ground between the detailed imagery of camera traps and the broad coverage of high-altitude surveys, enabling new behavioral and movement studies. Autonomous drones combined with robust detection models can transform wildlife monitoring from a static, pre-planned process to an adaptive system that responds to animal movements in real-time [26]. Deploying such systems involves important tradeoffs for ecologists. Lower flight altitudes provide better resolution but may increase animal disturbance, potentially altering natural behaviors [1, 37]. The balance between human and automated analysis also presents challenges [24]. Manual analysis by trained biologists remains the reliability standard but limits study scale, while automated detection can dramatically increase processing speed but requires domain-specific training to achieve necessary reliability. The dramatic improvements achieved through fine-tuning suggest that with appropriate dataset development, AI-assisted wildlife monitoring can achieve the reliability needed for scientific applications while significantly increasing wildlife studies’ temporal and spatial coverage. By contributing both the MMLA dataset and fine-tuned model weights to the research community, we hope to accelerate the development of robust computer vision solutions for wildlife conservation and establish a foundation for the next generation of intelligent, autonomous wildlife monitoring systems.
@@ 8 Acknowledgments
This work was supported by WildDroneEU (EU Horizon Europe MSCA grant 101071224), the U.S. National Science Foundation through the Imageomics Institute (NSF HDR Award 2118240), and the AI institute for Intelligent Cyberinfrastructure with Computational Learning in the Environment (ICICLE) (NSF grant OAC-2112606).
We thank Samuel Mutisya and William Njoroge at Ol Pejeta Conservancy for their support in facilitating the WildDrone 2025 Hackathon. We thank Jackson Miliko at the Mpala Research Centre for his support in facilitating the Kenyan Animal Behavior Recognition (KABR) project, whose data was used in this work. We thank Dan Beetem, The Wilds, and the Columbus Zoo & Aquarium for their support in facilitating data collection. All drone missions at The Wilds was conducted under the supervision of the Director of Animal Management, with permission from The Wilds Animal Care and Use Committee.